### v1

# Alignment and Adversarial Robustness: Are More Human-Like Models More Secure?

Abstract
Representational alignment refers to the extent to which a model’s internal representations mirror biological vision, offering insights into both neural similarity and functional correspondence. Recently, some more aligned models have demonstrated higher resiliency to adversarial examples, raising the question of whether more human-aligned models are inherently more secure. In this work, we conduct a large-scale empirical analysis to systematically investigate the relationship between representational alignment and adversarial robustness. We evaluate 118 models spanning diverse architectures and training paradigms, measuring their neural and behavioral alignment and engineering task performance across 106 benchmarks as well as their adversarial robustness via AutoAttack. Our findings reveal that while average alignment and robustness exhibit a weak overall correlation, specific alignment benchmarks serve as strong predictors of adversarial robustness, particularly those that measure selectivity towards texture or shape. These results suggest that different forms of alignment play distinct roles in model robustness, motivating further investigation into how alignment-driven approaches can be leveraged to build more secure and perceptually-grounded vision models.
1 Introduction
A longstanding goal in computer vision is to develop models that process images in a way that aligns with human perception. Representational alignment—how closely a model resembles biological vision—has been studied extensively with the goal of measuring, bridging, or increasing alignment in machine learning models Sucholutsky et al. (2024). Recent observations suggest that alignment may have implications beyond neuroscience: models that are more aligned with human perception have also exhibited increased robustness to adversarial examples—inputs with near-imperceptible perturbations that induce model misclassification— Dapello et al. (2020); Li et al. (2019), hinting at a deeper connection between alignment and security.
However, the relationship between representational alignment and adversarial robustness remains poorly understood. While the former seeks to align models with human cognition, adversarial examples in security highlight a fundamental misalignment: imperceptible perturbations can drastically degrade model accuracy while leaving human perception unaffected. Prior robustness techniques, such as adversarial training Madry et al. (2019), are computationally expensive and potentially vulnerable to new attack strategies. Meanwhile, alignment research has not systematically examined whether more human-aligned models are inherently more robust to adversarial attacks. A fundamental question remains: do these objectives complement each other, leading to better-aligned and more robust models, or do they introduce conflicting trade-offs?
In this work, we investigate the relationship between human alignment and robustness to adversarial examples in vision models through a diverse, large-scale empirical analysis. In our analysis, we study 118 models across different architectures and training schemes, measure their alignment across 106 different benchmarks on neural, behavioral, and engineering tasks via the BrainScore library Schrimpf et al. (2018). We then evaluate the adversarial robustness of these models using AutoAttack Croce & Hein (2020), a state-of-the-art ensemble attack.
In analyzing the correlations between model robustness and alignment, our findings reveal that while robustness is weakly correlated with vision alignment on average, certain alignment benchmarks serve as strong indicators of model robustness. Specifically, we find that the top six benchmarks that were most positively correlated with robust accuracy, even with strong perturbations, all measured a model’s selectivity towards texture. These results suggest that different forms of alignment play distinct roles in model robustness, motivating further investigation into how alignment-driven approaches can be leveraged to build more secure and perceptually-grounded vision models.
2 Background
Representational Alignment. Representational alignment studies the extent to which internal representations of machine learning models correspond to human cognitive processes. Early studies found that deep neural networks (DNNs) trained on large-scale image datasets develop hierarchical feature representations similar to those observed in the primate ventral stream, particularly in high-level visual areas like the inferior temporal (IT) cortex Yamins et al. (2013); Schrimpf et al. (2018). This led to efforts to quantify the alignment between artificial and biological vision, using techniques such as Representational Similarity Analysis (RSA) Kriegeskorte et al. (2008) and Centered Kernel Alignment (CKA) Kornblith et al. (2019). Current research in the area primarily focuses on measuring, bridging, and increasing both neural and behavioral alignment.
To improve alignment, researchers have proposed strategies that incorporate cognitive constraints or psychological priors into model architectures Dapello et al. (2020). Supervised fine-tuning with human-annotated datasets Dosovitskiy et al. (2021) ensures that learned representations align more closely with human-understandable features. Furthermore, novel techniques Muttenthaler et al. (2023); Li et al. (2019); Cheng et al. (2024) have been developed to encourage similarity between model activations and human neural responses as recorded through fMRI and EEG experiments. In this study, we use a comprehensive set of neural, behavioral, and engineering alignment metrics to quantify representational alignment.
Adversarial Examples. Although machine learning models have shown strong capabilities in achieving high accuracy across various tasks Liu et al. (2022); Dosovitskiy et al. (2021); Krizhevsky et al. (2017); He et al. (2016), they remain vulnerable to adversarial examples Croce & Hein (2020); Madry et al. (2019); Carlini & Wagner (2017); Goodfellow et al. (2015); Sheatsley et al. (2023). Adversarial examples are specially crafted inputs that contain perturbations which are imperceptible to humans, yet significantly decrease model accuracy. In computer vision systems, there have been many studies on developing attack algorithms, such as FGSM Goodfellow et al. (2015), PGD Madry et al. (2019), and AutoAttack Croce & Hein (2020). These methods aim to maximize model’s loss subject to constraints of perturbations defined by certain $\ell_{p}$-norms as follows:
$x_{adv}=\operatorname*{arg\,max}_{\left\|\delta\right\|_{p}\leq\epsilon}L(x+\delta,y)$
where $x$ and $y$ represent the original image and its predicted label, respectively, $\delta$ is the perturbation to solve for, and $L$ is the model’s loss function. The perturbation constraint $\epsilon$ is measured through an $\ell_{p}$-norm—most commonly $\ell_{\infty}$. While many works have historically evaluated the robustness of their model through the PGD attack Madry et al. (2019), it has been shown that “robust” models can often suffer from gradient masking, causing gradient-based attacks like PGD to fail Athalye et al. (2018), and leading to a sense of overestimated robustness. To overcome this, multiple attacks, including both white- and black-box attacks should be used Carlini et al. (2019). Thus, the AutoAttack ensemble Croce & Hein (2020) has become the de-facto standard for evaluating robustness.
3 Methods
Alignment. To measure alignment and download candidate models, we leverage the BrainScore Schrimpf et al. (2018) library. BrainScore provides a standardized framework for evaluating model similarity to biological vision through a set of neural, behavioral, and engineering benchmarks, supplying 106 benchmarks in total. These benchmarks quantify how closely a model’s internal representations and outputs correspond to neurophysiological recordings, human psychophysical behavior, and performance on engineered vision tasks. Neural alignment is measured by comparing activations from DNNs to neural recordings from primate visual cortex regions (e.g., V1, V2, V4, and IT), using similarity metrics like Representational Similarity Analysis (RSA) Kriegeskorte et al. (2008). Behavioral alignment assesses whether models replicate human psychophysical responses in object recognition and perturbation tests, while engineering alignment evaluates model robustness to controlled distortions, such as contrast reductions, or performance on out of distribution data.
In total, the BrainScore library has documented benchmark scores for 434 models. Out of those, there are 197 models available in their registry (the remaining 237 models were either submitted privately or have been deprecated). From the 197 models in the registry, we removed an additional 72 models because either loading the model produced a ClientError due to a moved or removed model hosting location or the model was incompatible with ImageNet (either does not output 1000 classes or expects video streams). After this, we had to discard an additional 7 models, which represented all the VOne class models Dapello et al. (2020) because they were not able to run on AutoAttack due to gradient alteration or masking, suggesting that previous results finding that VOne models are more robust to adversarial examples could have been due to overestimated robustness and highlighting the importance of evaluating robustness under comprehensive attack strategies. After this filtering process, we were left with 118 models (see Appendix A) for our evaluation.
Robustness. To evaluate the robustness of our models, we use AutoAttack Croce & Hein (2020); Croce et al. (2021), which serves as the standard for evaluating the robustness of neural networks due to its strong attack performance and fully automated parameter-free design. AutoAttack contains 4 attacks: APGD-CE, APGD-DLR, FAB, and Square Attack. By evaluating on AutoAttack, we are not only evaluating on the most performant attacks, but also integrating in both white-box attacks and black-box attacks which has been recommended in previous works to combat reporting overestimated robustness due to gradient masking or obfuscation Carlini & Wagner (2017).
To better understand how the relationship between adversarial robustness and alignment changes as attacks change, we evaluate the $\ell_{\infty}$ robustness of our models at three different epsilon levels: $\epsilon=\{\frac{0.25}{255},\frac{0.5}{255},\frac{1}{255}\}$ to represent adversaries at different capability levels and small, medium, and large image distortion levels. While these values are typically lower than what would be benchmarked on platforms such as RobustBench Croce et al. (2021), we choose these values with the goal of having a wide distribution of robust accuracies to identify separability between models, rather than the goal of bringing the model down to 0% accuracy as what is typically done.
4 Results
In this work, we hypothesize that there is a relationship between model robustness and alignment, due to the inherent similarity of the goals in each of these spaces. Here, we focus on answering the question are more aligned machine learning models more robust to adversarial examples?
(a) Neural Alignment
(b) Behavioral Alignment
(c) Engineering Task Performance
Figure 1: Average vision alignment score vs robust accuracy on neural, behavioral, and engineering benchmarks.
To facilitate our experiments, we use the BrainScore library v2.2.4 to measure alignment Schrimpf et al. (2018) and load models. Details on models evaluated can be found in Section 3. Once these models have been loaded and their alignment has been measured across the 106 alignment benchmarks, we evaluate their robustness using AutoAttack Croce & Hein (2020) from the TorchAttacks Kim (2021) library v3.5.1. The ImageNet Russakovsky et al. (2015) validation set is used for clean inputs to the model and serves as the starting point to generate adversarial examples. All experiments are run across 12 A100 GPUs with 40 GB of VRAM and CUDA version 11.1 or greater.
4.1 Average Alignment
We first investigate how well different classes of alignment predict the robustness of a model. Here, we study neural alignment, behavioral alignment, and engineering task performance. For each of these classes, we take the average score across all the benchmarks, giving us a single score for each model in the class. While many works have typically studied average vision alignment overall (i.e., the average of all the benchmarks across all classes), it has been shown that this can overemphasize behavioral alignment at the cost of neural alignment Ahlert et al. (2024). For each model, we then compute its robust accuracy against AutoAttack at 3 different values of epsilon $\epsilon=\{0.001,0.00196,0.00392\}$, which corresponds to $\{\frac{0.25}{255},\frac{0.5}{255},$ and $\frac{1}{255}\}$, respectively.
In Figure 1, we analyze the average score for neural alignment, behavioral alignment, and engineering task performance on the x-axis and the robust accuracy on the y-axis. Each dot represents a model, and the 3 colors correspond to the model’s robust accuracy at 3 different epsilon values. We compute the line-fit of the data at each epsilon value and report the statistical significance.
We find statistically significant correlations at: all $\epsilon$ values for neural alignment (explaining up to 13% of variance), $\epsilon=0.001$ for behavioral alignment (7.1% of variance), and at the two lowest $\epsilon$ values for engineering task performance (up to 27% of variance). Overall, the relatively low $R^{2}$ values, coupled with the difficulty of getting statistically significant correlations at higher epsilon values, suggests that average alignment scores are, at best, a weak indicator of robust accuracy.
4.2 Individual Benchmarks
Motivated by the previous experiment where we find that average alignment is weakly correlated with robust accuracy, we hypothesize this counter-intuitive result occurs because averaging scores across different benchmarks may obscure that some individual benchmarks are stronger predictors of robust accuracy than others. To further explore this hypothesis, we collect all models’ scores on individual benchmarks for the three classes (neural alignment, behavioral alignment, and engineering task performance) and compute the correlation between each of these scores and robust accuracy at our three different $\epsilon$ values. Figure 2 shows a heatmap of the 106 different benchmarks on the x-axis and robust accuracy at three different $\epsilon$ values on the y-axis. In each cell, we report the Pearson correlation coefficient between the selected benchmark score and robust accuracy across models.
Figure 2: Heatmap of each of the BrainScore benchmarks, ordered and separated (black bars) by area of alignment (neural, behavioral, engineering) vs the robust accuracy. Each cell represents the correlation between a benchmark across models and the robust accuracy for those models.
From this figure, we find multiple interesting trends. First, we see a wide range of correlations between different benchmarks, confirming our hypothesis that not every current alignment metric is a good indicator of robust accuracy. Additionally, we sometimes see significant changes to the correlation of robust accuracy and a benchmark as the $\epsilon$ value increases (and thus becomes a stronger attack). These changes appear to cluster by class of alignment. Roughly speaking, the neural alignment benchmarks (shown from the first to second black bar) tend to have more stable (and more positive) correlations as $\epsilon$ increases. The behavioral benchmarks (shown from the second to third black bars) tend to be, surprisingly, often negatively correlated with robust accuracy at mid and high $\epsilon$ values, and the correlation mostly decreases as $\epsilon$ increases. Finally, engineering task performance (shown from the third black bar to the end of the figure) seems highly dependent on the task, with benchmarks in this category having correlations at both ends of the spectrum.
Interestingly, we find some trends in the benchmarks that strongly correlate with robust accuracy. Most notably, many of the benchmarks that exhibit strong positive correlations with robust accuracy even at high values of epsilon tend to measure a models bias toward texture to some degree. In the neural category, we found strong postive correlations in FreemanZiemba2013.V1-pls and FreemanZiemba2013.V2-pls from Freeman et al. (2013), which measures neural responses in V1 and V2 to naturalistic texture stimuli. In the engineering category, two sets of benchmarks stood out as having strong correlations with robust accuracy. First is the Geirhos2021cueconflict-top1 benchmark from Geirhos et al. (2019), which measures the probability of a model classifying an object using shape information rather than texture via texture-shape conflicted images. The other is the set of benchmarks from Hermann et al. (2020): Hermann2020cueconflict-shape_bias and Hermann2020cueconflict-shape_match, which similarly measures the probability of a model classifying an object using shape information and the percentage of the times the model classifies according to the shape class, rather than texture or other classes.
5 Related Work
There has been substantial progress on bridging the representational differences between humans and machine learning models over the last few years. Geirhos et al. (2021) shows many of the high-performance models match or in many cases exceed human feedforward performance on most of the OOD datasets studied. New models have also been introduced to promote both alignment and robustness. For example, Dapello et al. (2020) designs a new block for CNNs called the VOne block, which simulates V1 area processing. This work found that incorporating the VOne block into ResNet models increased robustness to both white box adversarial examples and common corruptions without sacrificing clean performance on ImageNet. Li et al. (2019) introduced a technique for regularizing machine learning models based on human neural readings and found that the resultant regularized models were more robust and human-aligned.
Subramanian et al. (2023) shows that the property difference of the spatial frequency channel between humans and neural networks explains both shape bias and adversarial robustness of networks. Models with higher levels of human alignment have also been shown to be more robust to distribution shifts and ImageNet-A data Sucholutsky & Griffiths (2023). Additionally, it has been shown that models tend to prioritize texture information over shape information Geirhos et al. (2019); Hermann et al. (2020) and that this bias extends to real-data decisions and is one of the major causes for vulnerability to natural adversarial samples Hoak et al. (2024); Hoak & McDaniel (2024).
6 Conclusions
In this work, we find that, perhaps surprisingly, representational alignment and adversarial robustness in vision systems are not always correlated. However, we do observe that certain individual benchmarks serve as strong indicators of robust accuracy, particularly those that assess a model’s preference for texture information over shape. From this, we hope to encourage future work to leverage insights found in both areas to build more secure and aligned vision systems.
Acknowledgments
This material is based upon work supported by, or in part by, the National Science Foundation under Grant No. CNS 2343611, and by the Combat Capabilities Development Command Army Research Office under Grant No. W911NF-21-1-0317 (ARO MURI). Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the U.S. Government, or the Department of Defense. The U.S. Government is authorized to reproduce and distribute reprints for government purposes notwithstanding any copyright notation hereon.
Appendix A Appendix
Evaluated Models
grcnn vgg 19 vgg 16
alexnet CORnet-S ReAlnet10
ReAlnet09 ReAlnet08 ReAlnet07
ReAlnet05 ReAlnet04 ReAlnet03
ReAlnet02 grcnn 109 ReAlnet01
BiT-S-R50x1 AlexNet SIN BiT-S-R50x3
BiT-S-R101x1 densenet-121 densenet-169
resnet50-sup resnet50-SIN inception v1
nasnet large BiT-S-R152x2 densenet-201
BiT-S-R152x4 inception v4 inception v3
BiT-S-R101x3 nasnet mobile artResNet18 1
regnet y 400mf antialiased-r50 resnet50-barlow
imagenet l2 3 0 efficientnet b0 resnet50-SIN IN
effnetb1 272x240 resnet-50-robust Res2Net50 26w 4s
resnet50 tutorial convnext tiny sup resnet50-SIN IN IN
shufflenet v2 x1 0 AT efficientnet-b2 tv efficientnet-b1
resnet SIN IN FT IN resnet50-vicregl0p9 ViT L 32 imagenet1k
alexnet ks torevert antialias-resnet152 resnet50-vicregl0p75
resnext101 32x8d wsl resnet-152 v2 pytorch resnext101 32x16d wsl
resnext101 32x32d wsl resnext101 32x48d wsl resnet50 imagenet full
resnet34 imagenet full resnet18 imagenet full focalnet tiny lrf in1k
resnet50 robust l2 eps1 AdvProp efficientnet-b4 resnet50 robust l2 eps3
AdvProp efficientnet-b6 resnet152 imagenet full resnet101 imagenet full
AdvProp efficientnet-b8 AdvProp efficientnet-b7 AdvProp efficientnet-b2
convnext femto ols:d1 in1k antialiased-rnext101 32x8d resnet50 imagenet 1 seed-0
resnet50 imagenet 10 seed-0 cv 18 dagger 408 pretrained convnext tiny:in12k ft in1k
deit small imagenet 1 seed-0 resnet50 imagenet 100 seed-0 efficientnet b0 imagenet full
efficientnet b2 imagenet full efficientnet b1 imagenet full deit small imagenet 10 seed-0
deit base imagenet full seed-0 deit small imagenet 100 seed-0 deit large imagenet full seed-0
convnext large:fb in22k ft in1k deit small imagenet full seed-0 convnext small imagenet 1 seed-0
convnext xlarge:fb in22k ft in1k convnext small imagenet 10 seed-0 convnext base imagenet full seed-0
convnext tiny imagenet full seed-0 convnext small imagenet 100 seed-0 convnext small imagenet full seed-0
convnext large imagenet full seed-0 vit base patch16 clip 224:openai ft in1k vit large patch14 clip 224:openai ft in1k
vit large patch14 clip 224:laion2b ft in1k convnext xxlarge:clip laion2b soup ft in1k vit tiny r s16 p8 384:augreg in21k ft in1k
vit large patch14 clip 336:laion2b ft in1k vit relpos base patch16 clsgap 224:sw in1k effnetb1 cutmix augmix sam e1 5avg 424x377
vit relpos base patch32 plus rpn 256:sw in1k convnext base:clip laiona augreg ft in1k 384 vit base patch16 clip 224:openai ft in12k in1k
swin small patch4 window7 224:ms in22k ft in1k vit huge patch14 clip 224:laion2b ft in12k in1k vit large patch14 clip 224:openai ft in12k in1k
vit large patch14 clip 336:openai ft in12k in1k vit huge patch14 clip 336:laion2b ft in12k in1k vit large patch14 clip 224:laion2b ft in12k in1k
convnext large mlp:clip laion2b augreg ft in1k 384 resnet50 finetune cutmix AVGe2e3 robust linf8255 e0 247x234 effnetb1 cutmixpatch augmix robust32 avge4e7 manylayers 324x288
effnetb1 cutmixpatch SAM robust32 avge6e8e9e10 manylayers 324x288